# Intervening in Black Box: Concept Bottleneck Model for Enhancing Human Neural Network Mutual Understanding

### v1

@@ 1 Introduction
Beyond CBM-based approaches, numerous studies employ interpretable surrogate structures and human-understandable concept space to explain black-box behavior without compromising performance [11, 27]. While these methods enable transparent reasoning, answering questions like “Why is it/is it not”, they focus on explanation rather than intervention. Beyond CBM [24] addresses this limitation by using probe functions to map intermediate black-box representations to a human-understandable concept space. This approach identifies biases and aligns reasoning pathways by modifying the concept distribution based on prior knowledge. However, its interventions are restricted to the sample level, without directly modifying black-box parameters, and requires additional data to train the probe function for concept space projection.
@@ 2.2 Model Intervention
Model intervention encompasses techniques that go beyond identifying biases or errors in black-box models, allowing for direct adjustments that improve both accuracy and interpretability. For example, Moayeri et al. [26] used attribution methods to identify spurious correlations in models trained on Hard-ImageNet and applied foreground constraints to refine model attention. Yet, attribution methods typically indicate only “where” the model focuses, without explaining “what” specific features are influential. As a result, explanations are often limited to the foreground of the object, preventing fine-grained attribution to critical features.
@@ 3.5 Concepts Intervention
These two errors are represented by concept scores and quantified using contribution matrices $S_{\text{nT}},S_{\text{pF}}\in\mathbb{R}^{N_{\Gamma}\times N_{c}}$, which capture the statistical impact of each concept on global model behavior.
To mitigate the influence of unfavorable concepts, we rank them based on $S_{\text{nT}}$ and $S_{\text{pF}}$ in descending order, removing the top $q$ most detrimental concepts for each confused class. The concept contribution matrices $S_{\text{nT}}$ and $S_{\text{pF}}$ are evaluated using global gradient-based concept intervention function [32, 4] that defined as
$G(W,P(x_{i}))=\frac{\partial P(x_{i})}{\partial W}\odot W,$ (6)
where $\odot$ denotes element-wise multiplication.
Given ${P}_{y_{i}}(x_{i}):=[P_{\text{cbm}}(x_{i})]_{1,y_{i}}$ as the ground-truth prediction, we define the index of maximum value in $P_{\text{cbm}}(x_{i})$ as $y^{*}_{i}:=\arg\max_{j}[P_{\text{cbm}}(x_{i})]_{1,j}$, with corresponding maximum prediction $P_{y^{*}_{i}}(x_{i})$, the concept contribution matrices $S_{\text{nT}}$ and $S_{\text{pF}}$ are defined as $\left[G(W,P_{y_{i}}(x_{i}))\right]_{y_{i}}$ denotes $y_{i}$-th row of matrix $G$. $S_{\text{nT}}(i)_{y_{i}}$ and $S_{\text{pF}}(i)_{y^{*}_{i}}$ are computed similarly, and $\mathbf{0}$ denotes the all-zero vector.
$S_{\text{nT}}(i)_{y_{i}}$ :=$:=\begin{cases}-S_{i}\times\left[G(W,P_{y_{i}}(x_{i}))\right]_{y_{i}},\quad&\text{ if }y_{i}\in\Gamma\\ \phantom{-}\mathbf{0},\quad&\text{ otherwise}\end{cases}$ (7a)
$S_{\text{pF}}(i)_{y^{*}_{i}}$ :=$:=\begin{cases}\phantom{-}S_{i}\times\left[G(W,P_{y^{*}_{i}}(x_{i}))\right]_{y^{*}_{i}},\quad&\text{ if }y^{*}_{i}\neq y_{i}\\ \phantom{-}\mathbf{0},\quad&\text{ otherwise}\end{cases}.\vspace{-0.3in}$ (7b)
We summarize the concept intervention procedure described above in a more structured form in Algorithm 1.
Algorithm 1 Concepts Intervention
0:  $D_{\text{val}}$, $W$, $E_{\text{img}}$, and $E_{\text{text}}$.
1:  Initialize $S_{\text{nT}},S_{\text{pF}}\leftarrow 0$.
2:  for $i=1,2,3,\dots,N_{\text{val}}$ do
3:     Compute $C^{v}(i)$, $S_{i}$, and $P_{\text{cbm}}(x_{i})$ using (1), (3.3), and (3), respectively.
4:     Update corresponding $y_{i}$- or $y^{*}_{i}$-th row of contribution matrices $S_{\text{nT}}(i)$ and $S_{\text{pF}}(i)$ as in (7a) and (7b), and accumulate $S_{\text{nT}}$ and $S_{\text{pF}}$ iteratively.
@@
11:  return  Intervention concepts $I_{c}\in\mathbb{R}^{N_{\Gamma}\times\overline{q}}\mbox{ with }\overline{q}\leq q$.
After identifying the corresponding CBM intervention concepts’ index $I_{c}\in\mathbb{R}^{N_{\Gamma}\times\overline{q}}$ on $D_{\text{val}}$, we refine concept weight matrix $W\in\mathbb{R}^{N_{\Gamma}\times N_{c}}$ by pruning intervention concepts for each confused class. Specifically, we set the weights of all intervention concepts to zero within their respective confused classes, i.e., for each concept indexed by $j\in\{1,\dots,N_{c}\}$ corresponding to a confused class $k\in\Gamma$, we enforce $W[k,j]=0$ and denote the updated concept weight matrix as $\overline{W}$ for CBM.
@@ 4.1 Experiments Setups
Figure 2: Comparative visualization of intervention concepts and black-box attributions before and after intervention. It shows the visual results of NFResNet50 on Flower-102. Through concept intervention, CBM-HNMU identify and remove $\overline{q}$ harmful concept ($I_{c}$) from confusing class (“sword lily”) based on $S_{\text{nT}}$ and $S_{\text{pF}}$. After intervention knowledge is transferred back through CBM, the black-box model corrects the misclassified “bougainvillea”. Analyzing the visual concept changes in the black-box model’s middle layer before and after intervention, we observe a decrease in the “total Sobol indices” of the Related Visual Concept, which is corresponding to “prominent spadix in the center of the flower”. Compared with the bougainvillea sample marked by green border, we find that the flower center features of both classes are similar, which likely contributed to the misclassification. The intervention reduces the model’s reliance on such subtle distinctions, leading to more accurate predictions. For additional visual explanations, see Appendix. A.
@@ 4.2 Experiments Results
Herein, we evaluate the effectiveness of CBM-HNMU across multiple datasets by comparing global accuracy and fine-grained class predictions before and after black-box intervention. In addition, we also comparatively analyze attribution changes of the black box before and after the intervention based on the visualization results. Combined with the intervention process of the natural language concept bottleneck, we obtained the black box reasoning basis and intervention feature attribution that are easier for human to understand. Furthermore, we examine the impact of CBM-based classification methods (CBM and Post-hoc CBM) and the influence of the number of confused classes ($N_{\Gamma}$) on intervention effectiveness. In the end, to further validate the selected intervention concepts align with human visual perception and aid in correcting classification errors, we conducted a user study.
Performance Comparison. As shown in Table 1, the three architectures in our experiments achieved global accuracy improvements across five datasets after intervention with CBM-HNMU. To further verify the effect of CBM-HNMU, we conduct the identical comparative test using another three architectures on Flower-102, FGVC-Aircraft, and CUB-200. The results are presented in Table 2. As a result, CBM-HNMU achieved the highest accuracy improvement with NFResNet50 on CIFAR-10, increasing from $80.92\%$ to $83.56\%$, a gain of $2.64\%$. Additionally, Vit_Small achieved the highest average improvement (AVG IMP) across the five datasets, with an increase of $1.03\%$.
Table 3 shows the number of corrected error samples and the coverage of confusion classes for different intervention models on $D_{\text{test}}$ of Flower-102. The notation “$-n$” (left of the bracket) indicates that $n$ samples are originally misclassified into a confusion class but are corrected after intervention. Conversely, “$+n$” (right of the bracket) represents $n$ samples that are initially misclassified into any class but are reassigned to a specific confusion class post-intervention. We observed that most corrected samples in $D_{\text{test}}$ are associated with intervened classes – either originally misclassified as one of the confused classes or having a true label within the confused set but predicted as another class. Additionally, the classification accuracy of most non-confused classes remains unchanged after intervention. It can be observed that all black-box models corrected generalization errors, and thus improve classification performance.
@@
Table 3: Statistics of fine-grained class interventions in Flower-102. The second column indicates correction details: $-n$ represents $n$ samples initially misclassified as a confused class that is corrected post-intervention, while $+n$ denotes $n$ samples originally misclassified but reassigned to the right class ($\Gamma$) after intervention. “Corrected” refers to the total number of error samples corrected by each model on $D_{\text{test}}$, and “Coverage” represents the number of corrected samples associated with the confused classes.
Intervention Visualization. Herein, We analyze how the intervened black box’s performance improvement relates to changes in natural language and visual-related concepts. Figure 2 presents comparative results illustrating the intervention process on Flower-102 dataset, highlighting visual concept explanations before and after intervention. Samples are randomly selected from a subset, where the black-box model’s original classification errors on the test set are corrected following intervention. Each corrected sample includes at least one pre- or post-correction class related to the confused classes, ensuring a clear visual link to the intervention concept (refer to “Coverage” in Table 3).
@@
Intervention Concept Visual Masking. To further investigate the relationship between natural language concepts removed during intervention and the erroneous visual dependencies of the original black box, we represent the corresponding vision-related intervention concepts as masks and apply them to misclassified samples. From hypothesis, masking visual features associated with intervening concepts should improve the black box’s classification performance on these erroneous samples. As in Figure 3, it shows that masking generally leads to error correction or an increase in the black box’s confidence in the correct class.
Figure 3: Intervention concept visual masking results across models and datasets. Different confusion classes are color-coded along with their corresponding visually related intervened concepts. Notably, in the third row, since intervention concept of “Common Tern” appears “gray”, a white mask is used to avoid impact.
@@
Finally, since deleting concepts alone may not correct errors stemming from concept bottleneck limitations, we introduce a concept replacement method using an additional concept search set (see Appendix. B).
@@
User Study. Our model reduces reliance on misleading concepts tied to fine-grained confusion, leading to minor changes in attribution maps but notable shifts in concept importance scores. To validate the selected intervention concepts align with human perception and aid in correcting classification errors, we conducted a user study on Flower-102, CUB-200, and FGVC-Aircraft. For each confused class in the test set, we clustered samples into ten groups and selected one representative per cluster as reference examples for participants. We recruited $30$ paid, non-expert volunteers to assess each intervention concept according to whether it matched true class (CT), false class (CF), or semantically rational (CR), as shown in Figure 5.
@@
The results show that the user’s confidence in the correlation (CT and CF) between the natural language concepts selected for intervention by CBM-HNMU and the visual concepts in the confusing class samples and the possible classification errors between confusing classes (CR) are generally higher than $0.5$, and the matching confidence for the true class and the false class is basically consistent. This further shows that the intervention concepts selected by the proposed method are generally the misleading concepts in the confusing classes and are consistent with human visual perception.
@@ Acknowledgments
This work was supported by grants from the Natural Science Foundation of Shanxi Province (2024JCJCQN-66), Science and Technology Commission of Shanghai Municipality (NO.24511106900), and is partially supported by the National Natural Science Foundation of China under grant Nos. 62072358 and 62072352.
@@
A. Visualizations and Explanations
B. Concept Replacement
C. Approximation and Intervention for CBM
D. Detail for Intervention Concept Visual Masking
E. Hardware and Software Settings
F. Parameter Settings
G. Discussion
H. Limitations
@@
The organization of supplementary material is as follows: Section A presents additional visualizations of intervention explanations. To address the concept bottleneck limitations, Section B introduces the concept replacement method based on additional concept search sets. Section C illustrates the optimization of predictions before and after intervention within an interpretable CBM framework, along with an analysis of its approximation to a black-box model for generalization. A comparison of the change in classification accuracy for the non-intervention class before and after the intervention is also provided here. Section D details an intervention-based concept masking experiment for vision-related tasks. Finally, Sections E through H provide further experimental details and discussions.
Figure I: More visualization of NFResNet50 on Flower-102.
Figure II: More visualization of NFResNet50 on FGVC-Aircraft.
Appendix A Visualization and Explanation
We begin with additional visualizations of intervention-based explanations, as shown in Figures IIX. These illustrations highlight the relationship between natural language intervention concepts and changes in black-box visual representations before and after intervention. We present results using CBM-NHMU on NFResNet50, BotNet26, and RexNet100, with interventions applied to Flower-102, CUB-200, and FGVC-Aircraft. Notably, the samples are randomly selected from a subset where the black-box model’s original classification errors on the test set are corrected after intervention. Each corrected sample includes at least one pre- or post-correction class associated with the confused classes, ensuring a clear visual link to the intervention concept (see “Coverage” in the manuscript).
Figure III: More visualization of NFResNet50 on CUB-200.
Figure IV: More visualization of BotNet26 on Flower-102.
In Figure I, the baseline of NFResNet50 misclassifies sword lily as bougainvillea. The visually related intervention concept: ”Prominent spadix in the center of the flower.” prompts us to delete the related concept about the prominence in the center of flower to correct this error. We can find that the third most important visual concept extracted by the original black box before intervention corresponds exactly to this description, and the importance score of this concept is $0.11$. After the intervention, the black box gives a similar conceptual explanation, but the difference is that the importance score of the intervention concept dropped significantly ($\downarrow 0.06\rightarrow 0.05$). Combining the misclassifies sample image, we can also see that the visual features corresponding to the intervention concepts are indeed easy to confuse the two classes.
In Figure II, NFResNet50 misclassifies BAE 146-300 as BAE 146-200. The visually related intervention concept: ”Blue with a grey and white stripe.” and ”Large logo on the tail.” prompts us to delete the related concept about the tail and the color-related feature of aircraft to correct this error. We can find that the second and the third most important visual concept extracted by the original black box before intervention corresponds exactly to the descriptions. After the intervention, we can find that the network relies less on the corresponding color-related feature and tail of the aircraft to classify real class. The concept importance scores of color-related features and aircraft tail dropped from $0.22$ to $0.18$ and $0.15$ to $0.13$, respectively.
Figure V: More visualization of BotNet26 on FGVC-Aircraft.
Figure VI: More visualization of BotNet26 on CUB-200.
In Figure III, NFResNet50 misclassifies shiny cowbird as American crow. The visually related intervention concept: ”Black webbed foot.” prompts us to delete the related concept about the foot of bird to correct this error. We can find that the second and the third most important visual concept extracted by the original black box before intervention corresponds exactly to this description. After the intervention, we can find that the network relies less on bird footsteps to classify real class.
In Figure IV, BotNet26 misclassifies bougainvillea as siam tulip. The visually related intervention concept: ”Very feminine shade of pink.” prompts us to delete the related concept about the pink color of flower to correct this error. We can find that the first most important visual concept extracted by the original black box before intervention corresponds exactly to this description, and the importance score of this concept is $0.80$. After the intervention, the black box gives a similar conceptual explanation, but the difference is that the importance score of the intervention concept dropped significantly ($\downarrow 0.30\rightarrow 0.50$).
In Figure V, BotNet26 misclassifies Model B200 as DC-3. The visually related intervention concept: ”Retractable tricycle landing gear.” and ”tricycle landing gear configuration with nose wheel steering.” prompts us to delete the related concept about the gear of aircraft to correct this error. We can find that the second most important visual concept extracted by the original black box before intervention corresponds exactly to this description. After the intervention, we can find that the network relies less on gear of the aircraft to classify real class and the original score of the corresponding visual concept dropped from $0.24$ to $0.15$.
In Figure VI, BotNet26 misclassifies common tern as forsters tern. The visually related intervention concept: ”A duck with a long, thin bill.” prompts us to delete the related concept about the bill of bird to correct this error. We can find that the second and the third most important visual concept extracted by the original black box before intervention includes this description. After the intervention, we can find that the network relies less on bird bill to classify real class and the original score of the second most important visual concept dropped from $0.21$ to $0.09$. The third most important visual concept even disappears.
Figure VII: More visualization of RexNet100 on Flower-102.
Figure VIII: More visualization of RexNet100 on FGVC-Aircraft.
Figure IX: More visualization of RexNet100 on CUB-200.
In Figure VII, RexNet100 misclassifies water lily as camellia. The visually related intervention concept: ”heart-shaped spathe.”, ”curved spadix.” and ”prominent spadix in the center of the flower.” prompts us to delete the related concept about the pistil of flower to correct this error. We can find that the second most important visual concept extracted by the original black box before intervention includes corresponding features. However, the stamen feature given in concept $2$ can easily be confused between the two classes. After the intervention, we clearly can find that the black box replaced concept $2$ with a more representative visual concept of the stamen to water lily.
In Figure VIII, RexNet100 misclassifies MD-87 as MD-80. The visually related intervention concept: ”Large t-shaped tail fin.”, ”white with blue and red stripes.”, ”large tail fin with the airline’s logo.”, etc. prompts us to delete the related concept about the tail with corresponding color-related and shape-related feature of aircraft to correct this error. We can find that the second most important visual concept extracted by the original black box before intervention corresponds exactly to this description. After the intervention, we can find that the network relies less on the tail of the aircraft to classify real class and the original score of the corresponding visual concept dropped from $0.45$ to $0.35$.
In Figure IX, RexNet100 misclassifies white necked raven as American crow. The visually related intervention concept: ”Head is grayish-brown.” and ”large, black bird with a long, thick beak.” prompts us to delete the related concept about the head and beak of bird to correct this error. We can find that the first most important visual concept extracted by the original black box before intervention corresponds exactly to this description. After the intervention, we can find that the network relies less on bird head to classify real class and the original score of the corresponding visual concept dropped from $0.37$ to $0.28$.
Appendix B Concept Replacement

### v2

@@ 1 Introduction
Beyond CBM-based approaches, numerous studies employ interpretable surrogate structures and human-understandable concept space to explain black-box behavior without compromising performance [11, 27]. While these methods enable transparent reasoning, answering questions like “Why is it/is it not”, they focus on explanation rather than intervention. Beyond CBM [24] addresses this limitation by using probe functions to map intermediate black-box representations to a human-understandable concept space. This approach identifies biases and aligns reasoning pathways by modifying the concept distribution based on prior knowledge. However, its interventions are restricted to the sample level, without directly modifying black-box parameters, and require additional data to train the probe function for concept space projection.
@@ 2.2 Model Intervention
Model intervention encompasses techniques that go beyond identifying biases or errors in black-box models, allowing for direct adjustments that improve both accuracy and interpretability. For example, Moayeri et al. [26] used attribution methods to identify spurious correlations in models trained on Hard-ImageNet and applied foreground constraints to refine model attention. Yet, attribution methods typically indicate only “where" the model focuses, without explaining “what" specific features are influential. As a result, explanations are often limited to the foreground of the object, preventing fine-grained attribution to critical features.
@@ 3.4 Local Approximation
These two errors are represented by concept scores and quantified using contribution matrices $S_{\text{nT}},S_{\text{pF}}\in\mathbb{R}^{N_{\Gamma}\times N_{c}}$, which capture the statistical impact of each concept on global model behavior for each confusion class.
To construct the concept contribution matrices $S_{\text{nT}}$ and $S_{\text{pF}}$, consider a sample $x_{i}\in D$ with ground-truth label $y_{i}$. Let $y^{*}_{i}:=\displaystyle\max_{j}[P_{\text{cbm}}(x_{i})]_{1,j}$ denotes predicted class under the concept bottleneck model and $W\in\mathbb{R}^{N_{\Gamma}\times N_{c}}$ the concept weight matrix, where each row vector $w_{k}\in\mathbb{R}^{1\times N_{c}}$ corresponds to confusion class $k$.
To assess the influence of concepts on model behavior, we define a gradient-based attribution function
$G(w_{k},P_{k}(x_{i})):=\frac{\partial P_{k}(x_{i})}{\partial w_{k}}\odot w_{k},$ (6)
where $\odot$ denotes element-wise multiplication and $P_{k}(x_{i})=[P_{\mathrm{cbm}}(x_{i})]_{1,k}$ is the predicted probability for class $k$. Based on $S_{i}\in\mathbb{R}^{1\times N_{c}}$ from Eq. (3.3), the per-sample contribution scores are defined as follows:
$S_{\text{nT}}(i)_{y_{i}}$ :=$:=\begin{cases}-S_{i}\odot G(w_{y_{i}},P_{y_{i}}(x_{i})),\;\;&\text{ if }y_{i}\in\Gamma\\ \phantom{-}\mathbf{0},\;\;&\text{ otherwise}\end{cases}$ (7a)
$S_{\text{pF}}(i)_{y^{*}_{i}}$ :=$:=\begin{cases}\phantom{-}S_{i}\odot G(w_{y_{i}^{*}},P_{y^{*}_{i}}(x_{i})),\;\;&\text{ if }y^{*}_{i}\neq y_{i}\\ \phantom{-}\mathbf{0},\;\;&\text{ otherwise}\end{cases}.$ (7b)
Here, $S_{\text{nT}}$ captures how the activated concepts suppress the true class, while $S_{\text{pF}}$ measures how they reinforce the incorrect predicted class. To mitigate the impact of unfavorable concepts, we accumulate the sample-wise scores over the validation set to form global contribution matrices. Concepts in each confused class are ranked by their total influence in $S_{\text{nT}}$ and $S_{\text{pF}}$, and the top $\overline{q}$ most detrimental concepts are selected for removal during intervention.
We summarize the concept intervention procedure in Algorithm 1, which identifies and prunes detrimental concepts based on their contribution to prediction errors.
Algorithm 1 Gradient-Based Concept Intervention
0:  Validation set $D_{\text{val}}$, concept weight matrix $W$, encoders $E_{\text{img}}$, $E_{\text{text}}$
1:  Initialize contribution matrices $S_{\text{nT}}\leftarrow 0$, $S_{\text{pF}}\leftarrow 0$
2:  for $i=1$ to $N_{\text{val}}$ do
3:     Compute $S_{i}$, $P_{\text{cbm}}(x_{i})$, and predicted class $y_{i}^{\star}$
4:     Update $S_{\text{nT}}(i)_{y_{i}}$ and $S_{\text{pF}}(i)_{y_{i}^{\star}}$ using Eqs. (7a) and (7b), then accumulate $S_{\text{nT}}$ and $S_{\text{pF}}$
@@
11:  return  Intervention indices $I_{c}\in\mathbb{R}^{N_{\Gamma}\times\bar{q}}$ with $\bar{q}\leq q$
After identifying the intervention index set $I_{c}$ from $D_{\text{val}}$, we update the concept weight matrix $W\in\mathbb{R}^{N_{\Gamma}\times N_{c}}$ by zeroing out the selected concept weights for each confused class. Specifically, for each $k\in\Gamma$ and each concept $j\in I_{c}^{k}$, we set $W[k,j]=0$. The resulting $\overline{W}$ defines the intervened concept weight matrix of the CBM.
@@ 4.1 Experiments Setups
Figure 2: Comparative visualization of intervention concepts and black-box attributions before and after intervention. It shows the visual results of NFResNet50 on Flower-102. Through concept intervention, CBM-HNMU identify and remove $\overline{q}$ harmful concept ($I_{c}$) from confusing class (“sword lily") based on $S_{\text{nT}}$ and $S_{\text{pF}}$. After intervention knowledge is transferred back through CBM, the black-box model corrects the misclassified “bougainvillea". Analyzing the visual concept changes in the black-box model’s middle layer before and after intervention, we observe a decrease in the “total Sobol indices" of the Related Visual Concept, which is corresponding to “prominent spadix in the center of the flower". Compared with the bougainvillea sample marked by green border, we find that the flower center features of both classes are similar, which likely contributed to the misclassification. The intervention reduces the model’s reliance on such subtle distinctions, leading to more accurate predictions. For additional visual explanations, see Appendix. H.
@@ 4.2 Experiments Results
We evaluate the improvements in both global accuracy and fine-grained class prediction, resulting from the CBM-HNMU intervention across multiple datasets. In addition, we examine how the visual concepts captured by black-box model evolve before and after the intervention, using visualization techniques and natural language concepts to interpret the process. An ablation study further illustrates how the quality of black-box approximation and the overall intervention effect vary with the number of intervention concepts. Furthermore, we examine the impact of CBM-based classification methods (CBM and Post-hoc CBM) and the influence of the number of confused classes ($N_{\Gamma}$) on intervention effectiveness. Finally, we conducted a user study to confirm that the chosen intervention concepts (i) align with human visual perception and (ii) help correct classification errors.
Performance Improvements. As shown in Table 1, the three architectures in our experiments achieved global accuracy improvements across five datasets after intervention with CBM-HNMU. To further verify the effect of CBM-HNMU, we conduct the identical comparative test using another three architectures on Flower-102, FGVC-Aircraft, and CUB-200. The results are presented in Table 2. As a result, CBM-HNMU achieved the highest accuracy improvement with NFResNet50 on CIFAR-10, increasing from $80.92\%$ to $83.56\%$, a gain of $2.64\%$. Additionally, Vit_Small achieved the highest average improvement (AVG IMP) across the five datasets, with an increase of $1.03\%$.
Table 3 shows the number of corrected samples and the coverage of confusion classes in different intervention models on $D_{\text{test}}$ of Flower-102. The notation “$-n$” indicates that $n$ samples originally misclassified into a confusion class are corrected after intervention. Conversely, “$+n$” denotes $n$ samples initially misclassified into any class that are reassigned to the confusion class after intervention. Most corrected samples in $D_{\text{test}}$ are related to the intervened classes, either because they were previously predicted as one of the confusion classes or because their true labels belong to the confusion set.
@@
Table 3: Statistics of fine-grained class interventions in Flower-102. “Corrected” refers to the total number of error samples corrected by each model on $D_{\text{test}}$, and “Coverage” represents the number of corrected samples associated with the confused classes.
Intervention Visualization. Herein, We analyze how the intervened black box’s performance improvement relates to changes in natural language and visual-related concepts. Figure 2 presents comparative results illustrating the intervention process on Flower-102 dataset, highlighting visual concept explanations before and after intervention. Samples are randomly selected from a subset, where the black-box model’s original classification errors on the test set are corrected following intervention. Each corrected sample includes at least one pre- or post-correction class related to the confused classes, ensuring a clear visual link to the intervention concept (refer to “Coverage" in Table 3).
@@
User Study. We tested whether the concepts suppressed by CBM-HNMU are the same cues that humans find misleading. The study spanned the Flower-102, CUB-200, and FGVC-Aircraft test sets. For each pair of confused class we clustered its images into ten groups and retained one exemplar per cluster. Thirty paid, non-expert volunteers examined each exemplar along with its intervention concept and scored the concept as (i) true-class cue (CT), (ii) false-class cue (CF), or (iii) a semantically related cue shared by both classes (CR). As shown in Fig. 4, the mean confidence for all three score exceeds $0.5$, and the CT and CF scores are nearly identical. This parity, together with the high CR confidence, indicates that CBM-HNMU targets exactly the visual hints that mislead both the network and human observers, validating the choice of intervention concepts.
@@ Acknowledgments
This work was supported by grants from the Natural Science Foundation of Shanxi Province (2024JCJCQN-66), Science and Technology Commission of Shanghai Municipality (NO.24511106900), Key R&D Program of Zhejiang (2024SSYS0091) and is partially supported by the National Natural Science Foundation of China under grant Nos. 62072358 and 62072352.
@@
A. Concept Replacement
B. More Evaluation on Approximation and Intervention
C. Intervention Concept Visual Masking
D. Hardware and Software Settings
E. Parameter Settings
F. Discussion
G. Limitations
H. Visualizations and Explanations
@@
The organization of supplementary material is as follows: Section A introduces the concept replacement method based on additional concept search sets. Section B illustrates the optimization of predictions before and after intervention within an interpretable CBM, along with an analysis of its approximation to a black-box model for generalization. Also, the change in black-box classification accuracy for non-intervened classes before and after the intervention is reported. Section C details an intervention-based concept masking experiment for vision-related tasks. Sections D through G provide further experimental details and discussions. Finally, Section H presents additional visualizations of intervention explanations. To address the concept bottleneck limitations.
Appendix A Concept Replacement